# In-context learning of evolving data streams with tabular foundational models

### v1

@@ 3 In-context stream learning
This sections offers insights into the drivers behind the ICL ability, and how it addresses the core challenges of adaptive learning in dynamic environments. ICL requires a model to implicitly construct a map from in-context examples, of a new concept on which it has not been previously trained, to a predictor without any updates to the model’s parameters themselves. Indeed, it has been show empirically that transformers with fixed parameters can learn linear functions comparable to the optimal least squares estimator, by encoding smaller models in their activations, and updating these implicit models as new examples appear in the context on the fly [84]. Furthermore, it has been demonstrated that transformers implement more complex function classes implicitly, i.e., sparse linear functions, two-layer neural networks, and decision trees, with performance that matches or exceeds task-specific learning algorithms [85]. However, it remains unclear to what extent these models are able to learn new tasks from in-context examples alone as opposed to indexing into a vast set of known tasks from the training data. In this regard, it has been hypothesized that this emergent behavior of transformers is driven by both the distributional qualities of the training data itself [86] and an architectural bias where the internal structure of the model ends up performing mesa-optimization [87].
@@
Natural data hypothesis. The data hypothesis was inspired by the observation that many natural data sources differ from typical batch supervised datasets due to a few notable features: (1) natural data is temporally bursty, i.e. a given concept may have a distribution that is not uniform across time, instead tending to appear in clusters [88]; (2) the meaning of these concepts in natural data is often dynamic rather than fixed, usually in a context-dependent way; (3) the marginal distribution across entities is highly skewed, with large numbers of rarely occurring classes [86]. Indeed, these are the properties of data streams that motivate the development of incremental learning algorithms. While batch i.i.d. training typically consists of item classes that reoccur with uniform regularity, and input-output mappings that are fixed throughout training, data stream mining implies training on sequences of data with drifting concepts, akin to having many labels per item [89], a few highly reoccurring concepts [46], and a large number of outliers that reoccur much more rarely, thus excluded from the concept forming process due to their disproportionately less likelihood to occur multiple times within a given context window [90].
Figure 2: ICL as a mesa-optimization capability
@@
Mesa-optimization hypothesis. The algorithm hypothesis identifies this emergent behavior as stemming from an architectural bias where the internal structure of transformers ends up performing mesa-optimization [87], which doesn’t happen for other recurrent models, like LSTMs and RNNs, even if matched on number of parameters [86]. Mesa-optimization occurs when a base optimizer is searching for algorithms to solve some problem and finds a model that is itself a mesa-optimizer [91]. Unlike the base objective, the mesa-objective is not specified directly by the programmers. Instead, a learned process running within the forward pass constructs an internal mesa-objective, and its corresponding solution through mesa-optimization, because it is instrumentally useful for solving the given task. While meta means to go above, thus stepping outside to a higher logical level, mesa means going to the other direction, to go down into the content narrowing the details into smaller and smaller pieces [92]. Many reasons have been identified for this phenomenon. Firstly, the base optimizer’s incentive to find a compressed policy creates a bias towards simple solutions with low description length. Moreover, the transformer’s statefulness favors mesa-optimization by decreasing the implicit penalty on time complexity imposed by enforcing all computations to end when the learned algorithm produces its output. And, the ability of the learned algorithm to save and refer back to previous state enables the caching of intermediate results. These factors, when combined with heavy pre-training on diverse tasks that are likely to be completely novel, incentivize the base optimizer to expend optimization power in learning how to adjust to a new task immediately in order to perform consistently in an diverse environment. Figure 2 illustrates this ability.
@@
Implicit meta-learning. While breaking things down into bits is a mesa process, the model always returns to rebuild and re-map the actual details. In going mesa, all generalizations are deconstructed which allows to reconstruct the input-output mapping in a more resourceful manner. In this regard, training a model to perform ICL can also be viewed as an instance of the learning-to-learn or meta-learning paradigm, with learning that takes place on various time scales, fast and slow, comprising both mesa- and meta-processes. Examples include approximating an implicit model selection procedure competitive with empirical risk minimization [93], a gradient-based few-shot learning within its forward pass [94] or a kernel-based sequence model [95]. However, in these cases, researchers explicitly designed the meta-training regime to incentivize ICL, whereas in the case of transformers, the capacity for ICL is emergent. Neither the model’s transformer architecture nor its learning objective are explicitly designed with ICL in mind.
Figure 3: ICL as bayesian inference
@@
Bayesian perspective. While the distinction between mesa- and meta-learning offer a perspective on how systems learn and adapt, a Bayesian theory can help to understand and quantify the uncertainty in the process of learning itself. In this regard, ICL can be viewed as approximate Bayesian inference, whether implicitly [96] or explicitly [97]. It has be shown that a Transformer directly fits the posterior predictive distribution (PPD), with a hypothesis latent concept $\phi$ which could parametrize the transitions of a Hidden Markov model [98], or a space of structural causal models. In this framework, the prior defines a space of hypotheses $\Phi$ on the relationship of a set of inputs $x$ to the output labels $y$. Each hypothesis $\phi\in\Phi$ can be seen as a mechanism that generates a data distribution from which one can draw context data with observed labels ${x_{1},y_{1},\ldots,x_{n},y_{n}}$ and an unlabeled query sample $x_{query}$, as illustrated in Figure 3.
@@
The PPD specifies the distribution of the label $p(y_{query}|x_{query},x_{1},y_{1},\ldots,x_{n},y_{n})$ and can be obtained by integration over the space of hypotheses $\Phi$, where the weight of a hypothesis $\phi\in\Phi$ is determined by the posterior probability $p(\phi|x_{1},y_{1},\ldots,x_{n},y_{n})$ given the context data. Considering the unnormalized posterior, which corresponds to the prior probability $p(\phi)$ and the likelihood $p(x_{1},y_{1},\ldots,x_{n},y_{n}|\phi)$ of the training data given $\phi$, we have:

### v2

@@ 2. In-context stream learning
ICL requires a model to implicitly construct a map from in-context examples to a predictor of a new concept on which it has not been previously trained, without any updates to the model’s parameters themselves. Indeed, it has been show empirically that transformers with fixed parameters can learn linear functions comparable to the optimal least squares estimator, by encoding smaller models in their activations, and updating these implicit models as new examples appear in the context on the fly (von2022what). Furthermore, it has been demonstrated that transformers implement more complex function classes implicitly, i.e., sparse linear functions, two-layer neural networks, and decision trees, with performance that matches or exceeds task-specific learning algorithms (min2022rethinking). However, it remains unclear to what extent these models are able to learn new tasks from in-context examples alone as opposed to indexing into a vast set of known tasks from the training data. In this regard, it has been hypothesized that this emergent behavior of transformers is driven by both the distributional qualities of the training data itself (chan2022data) and an architectural bias where the internal structure of the model ends up performing mesa-optimization (bai2022uncovering).
@@
Natural data hypothesis. The data hypothesis was inspired by the observation that many natural data sources differ from typical batch supervised datasets due to a few notable features: (1) natural data is temporally bursty, i.e. a given concept may have a distribution that is not uniform across time, instead tending to appear in clusters (altmann2009beyond); (2) the meaning of these concepts in natural data is often dynamic rather than fixed, usually in a context-dependent way; (3) the marginal distribution across entities is highly skewed, with large numbers of rarely occurring classes (chan2022data). Indeed, these are the properties of data streams that motivate the development of incremental learning algorithms. While batch i.i.d. training typically consists of item classes that reoccur with uniform regularity, and input-output mappings that are fixed throughout training, data stream mining implies training on sequences of data with drifting concepts, akin to having many labels per item (read2023multi), a few highly reoccurring concepts (anderson2016recurring), and a large number of outliers that reoccur much more rarely, thus excluded from the concept forming process due to their disproportionately less likelihood to occur multiple times within a given context window (de2016minas).
@@
Mesa-optimization hypothesis. The algorithm hypothesis identifies this emergent behavior as stemming from an architectural bias where the internal structure of transformers ends up performing mesa-optimization (bai2022uncovering), which doesn’t happen for other recurrent models, like LSTMs and RNNs, even if matched on number of parameters (chan2022data). Mesa-optimization occurs when a base optimizer is searching for algorithms to solve some problem and finds a model that is itself a mesa-optimizer (hubinger2019risks). Unlike the base objective, the mesa-objective is not specified directly by the programmers. Instead, a learned process running within the forward pass constructs an internal mesa-objective, and its corresponding solution through mesa-optimization, because it is instrumentally useful for solving the given task. While meta means to go above, thus stepping outside to a higher logical level, mesa means going to the other direction, to go down into the content narrowing the details into smaller and smaller pieces (cheal2021what). Many reasons have been identified for this phenomenon. Firstly, the base optimizer’s incentive to find a compressed policy creates a bias towards simple solutions with low description length. Moreover, the transformer’s statefulness favors mesa-optimization by decreasing the implicit penalty on time complexity imposed by enforcing all computations to end when the learned algorithm produces its output. And, the ability of the learned algorithm to save and refer back to previous state enables the caching of intermediate results. These factors, when combined with heavy pre-training on diverse tasks that are likely to be completely novel, incentivize the base optimizer to expend optimization power in learning how to adjust to a new task immediately in order to perform consistently in an diverse environment. Figure 2 illustrates this ability.
Figure 2. ICL as a mesa-optimization capability
@@
Implicit meta-learning. While breaking things down into bits is a mesa process, the model always returns to rebuild and re-map the actual details. In going mesa, all generalizations are deconstructed which allows to reconstruct the input-output mapping in a more resourceful manner. In this regard, training a model to perform ICL can also be viewed as an instance of the learning-to-learn or meta-learning paradigm, with learning that takes place on various time scales, fast and slow, comprising both mesa- and meta-processes. Examples include approximating an implicit model selection procedure competitive with empirical risk minimization (agrawal2022transformers), a gradient-based few-shot learning within its forward pass (von2022transformers) or a kernel-based sequence model (mccoy2022recasting). However, in these cases, researchers explicitly designed the meta-training regime to incentivize ICL, whereas in the case of transformers, the capacity for ICL is emergent. Neither the model’s transformer architecture nor its learning objective are explicitly designed with ICL in mind.
3. Large tabular model
Since the pre-training stage does not represent the added value of this work, we rely on the popular TabPFN transformer-based model (hollmann2022tabpfn; hollmann2025accurate), relying on its original dataset generation and architecture, only performing minor modifications on the inference time procedure. At the core of TabPFN is the self-attention mechanism, interchanging information among different positions within a sequence (vaswani2017attention). In self-attention, each element in the input sequence of length $T$ is linearly projected to produce $d$-dimensional queries, keys, and values, each represented by $Q,K,V\in\mathbb{R}^{T\times d}$. The attention matrix is computed as $\text{softmax}(QK^{T}/\sqrt{d})$ where softmax is applied row-wise. It is then multiplied to the value matrix to produce the output of the layer $O=\text{softmax}(QK^{T}/\sqrt{d})V$. If self-attention is computed in this manner, it is more precisely called bi-directional self-attention since every token can attend to all the other tokens in both directions. However, TabPFN relies on causal attention layers typical of a standard decoder-only Transformer architecture (liu2018generating), making it possible to process a streaming sequence where the tokens become available one at a time (mccoy2022recasting). The complete architecture is given in Figure 3.
Figure 3. Architecture and training procedure of the foundational model, where $\otimes$ represents features and float targets being separately encoded using a linear layer, and the embedding is pushed through a decoder-only transformer model
Architecture. Each instance’s feature vector is treated as a set-valued input to exploit permutation invariance and represented as a token with dimension $d_{\text{token}}$, allowing token representations to attend to each other for a fixed number of features $d_{f}$. Firstly, the creation of in-context tokens $H\in\mathbb{R}^{|S|\times d_{\text{token}}}$ is perfomed by embedding the features and classes linearly using weights $W_{x}\in\mathbb{R}^{d_{f}\times d_{\text{token}}}$ and $w_{y}\in\mathbb{R}^{d_{\text{token}}}$ respectively. Then, the embedded tokens are pushed through the standard decoder-only Transformer (liu2018generating). Each layer in the transformer autoregressively maps the matrix $H^{l}$ to a sequence $H^{l+1}$, processing each collumn $h_{i}^{l}$ of $H^{l}$ in parallel. Each layer first computes a self attention and then applies a feedforward transformation. However, with a special attention mask, where in-context tokens only attend to each other, with no attention to the single query instance, as illustrated in Figure 4. Instead, the target to the held-out query instance is used as a label, minimizing the cross-entropy loss, which approximates the true Bayesian posterior predictive distribution (muller2021transformers). This procedure results in a computational cost of a forward pass scaling as $O(|S|^{2})$.
Figure 4. Special attention mask
Learning. To prior-fit the network as part of the algorithm development, TabPFN repeatedly samples synthetic realistic datasets. These are generated by carefully crafting correlations blockwise, imposing causal relationships between features, and tuning the magnitude of weights for variety in the feature importance, using Bayesian Neural Networks and Structural Causal Models with random characteristics and randomly initialized weights (hollmann2022tabpfn; von2022what). In this regard, to better understand and quantify the uncertainty in the process of learning a LTM, one can take both a bayesian (xie2022explanation; garg2022transformers) and a frequentistic perspective (nagler2023statistical).
@@ 3. Large tabular model
Bayesian perspective. In-context learning can be viewed as approximate Bayesian inference, whether implicitly (xie2022explanation) or explicitly (garg2022transformers). It has be shown that a Transformer directly fits the posterior predictive distribution (PPD), with a hypothesis latent concept $\phi$ which could parametrize the transitions of a Hidden Markov model (baum1966statistical), or a space of structural causal models. In this framework, the prior defines a space of hypotheses $\Phi$ on the relationship of a set of inputs $x$ to the output labels $y$. Each hypothesis $\phi\in\Phi$ can be seen as a mechanism that generates a data distribution from which one can draw context data with observed labels $x_{1},y_{1},\ldots,x_{n},y_{n}$}, denoted as $c_{txt}$}, and an unlabeled query sample $x_{q}$, as illustrated in Figure 5.
Figure 5. ICL as Bayesian inference
@@
The PPD specifies the distribution of the label $p(y_{q}|x_{q},c_{txt})$ and can be obtained by integration over the space of hypotheses $\Phi$, where the weight of a hypothesis $\phi\in\Phi$ is determined by the posterior probability $p(\phi|c_{txt})$ given the context data. Considering the unnormalized posterior, which corresponds to the prior probability $p(\phi)$ and the likelihood $p(c_{txt}|\phi)$ of the training data given $\phi$, we have:
(1) $p(y_{q}|c_{txt},x_{q})=\int_{\Phi}p(y_{q}|x_{q},\phi)p(c_{txt}|\phi)p(\phi)d\phi$
Frequentistic perspective. Despite contexts being sampled from a different distribution than the pretraining distribution, the asymptotic prediction error of in-context learning is optimal when the signal about the latent concept in each context example is larger than the error due to the distribution mismatch. As a more informative and representative input space is provided, the learning error decreases (xie2022explanation). In this regard, one can also take a frequentistic perspective (nagler2023statistical). From a predictor’s variance standpoint, being a pre-tuned, but untrained predictor with many hyperparameters and multi-head attention, a LTM has extremely high sensitivity to individual training samples, which translates in an increased ability to choose submodels and vanishing variance. From a bias standpoint, hyperparameters are pre-tuned to be optimal for a set of tasks defined by the prior. If this prior has large enough support and does not concentrate too much away from the true hypothesis, one can guarantee that the PPD converges to a close approximation of the true predictive distribution. Consequently, whether the LTM predictor can learn at inference time depends on its structural properties, with the optimal LTM approximation being characterized by a Kullback-Leibler criterion (nagler2023statistical). To allow for accurate approximation of the conditional class probabilities, one needs sufficiently complex LTM models and prior. Practically, a LTM is trained on simulated data sets. The larger these data sets are, the more complex the PPD is approximated. The training set size can therefore be understood as a regularizer on the expected complexity of the network. Thus, while LTM implementations ensure only vanishing predictor’s variance, the bias vanishes only if the provided data is somewhat localized around the test feature.
4. Context optimization
Building on the theoretical insights into how LTMs can perform in-context stream learning, we now turn to the challenge of regulating their behavior in non-stationary environments. Specifically, the objective of context optimization is to minimize bias in model predictions when faced with new data that drifts over time. This problem occurs in situations where the LTM $f_{\theta}$, is applied on-the-fly to a continuous stream of real-world data. The data stream is assumed to be partitioned sequentially into non-overlapping segments, called concepts, such as $C_{i}$ and $C_{i+1}$, each corresponding to distinct joint data distributions, $p_{i}(x,y)$ and $p_{i+1}(x,y)$, respectively. Within each concept, the data distribution is assumed static, however, as the data stream progresses, drift between consecutive concepts can introduce discrepancies between past and future data. For instance, if concept $C_{i}$ has a skewed label distribution $p_{i}(y)$, and the following concept $C_{i+1}$ has a uniform label distribution $p_{i+1}(y)$, then the predictive model must adapt to account for this shift in label distribution (nejjar2024context). The goal is to select an optimal localized context $l$, denoted by $(c_{txt},x_{q})$, which helps $f_{\theta}$ approximate the correct output $y_{q}$ even for cases in the tail regions of these drifting distributions. In this case, the model’s expected prediction error for a new sample $(x_{q},y_{q})$ is defined by:
(2) $\begin{split}&\mathbb{E}_{l}[(f_{\theta}(x_{q}\mid c_{txt})-\mathbb{E}[y_{q}\mid x_{q}])^{2}]\\ &=\mathbb{E}_{l}[((f_{\theta}(x_{q}\mid c_{txt})-\mathbb{E}_{l}[f_{\theta}(x_{q}\mid c_{txt})])\\ &+(\mathbb{E}_{l}[f_{\theta}(x_{q}\mid c_{txt})]-\mathbb{E}[y_{q}\mid x_{q}]))^{2}]\\ &=\text{Var}[f_{\theta}(x_{q}\mid c_{txt})]+\left(\text{Bias}^{2}[f_{\theta}(x_{q}\mid c_{txt})]\right)+\sigma^{2}\end{split}$
LTM’s variance decreases with more in-context examples. Assume $f_{\theta}$ to be $c$-Lipschitz continuous with constant $c=(c_{1},\ldots,c_{n})$ where each $c_{i}$ scales as $\delta i^{-\alpha}$, $\delta$ is a positive constant, and $\alpha>0.5$. Then, for any two independently sampled contexts $(c_{txt},x_{q})$ and $(c_{txt}^{\prime},x_{q})$ from an unbounded stream $D$, the following inequality $|f_{\theta}(x_{q}\mid(c_{txt}))-f_{\theta}(x_{q}\mid(c_{txt}^{\prime}))|\leq\sum_{i=1}^{n}c_{i}\mathbf{1}_{\{x_{i}\neq x^{\prime}_{i}\}}$ holds (garg2022transformers). Applying McDiarmid’s Inequality (mcdiarmid1989method), for any $t>0$, the tail probability bound is:
(3) $Pr(|f_{\theta}(x_{q}\mid c_{txt})-\mathbb{E}_{l}[f_{\theta}(x_{q}\mid c_{txt})]|\geq t)\leq 2\exp\left(-\frac{2t^{2}}{|c|^{2}_{2}}\right)$
where $|c|^{2}_{2}=\sum_{i=1}^{\infty}(\delta i^{-\alpha})^{2}$, converging due to $\alpha>0.5$. By the Borel-Cantelli lemma, $f_{\theta}(x_{q}\mid c_{txt})$ converges almost surely to the expected prediction $\mathbb{E_{s}}[f_{\theta}(x_{q}\mid c_{txt})]$ as $n\rightarrow\infty$ (nejjar2024context). This setup implies that as more samples are added to the prompt, the model’s sensitivity to small input changes diminishes, reducing error volatility.
LTM’s biases inherited from in-context examples. Assume $f_{\theta}$ mimics gradient descent steps on examples $(c_{txt},x_{q})$, the model’s average prediction over many samples $\mathbb{E}[f_{\theta}(x_{q}\mid c_{txt})]$ almost surely converges to the true conditional expectation $\mathbb{E}[y_{q}\mid x_{q}]$ as $n\to\infty$, provided there is no drift in the data distribution (von2022transformers). However, if there is shift in the distribution, $(c_{txt},x_{q})$ may become misaligned with the current distribution, leading to biased predictions that reflect outdated patterns. This creates a tension between retaining examples from old concepts $\{C_{j}\}_{j=1}^{i-1}$ just for the global representation of all classes $y\in\mathcal{Y}$, and adapting to transient local patterns representing the most up to date concept $C_{i}$, at the cost of not representing all classes.
Dual-memory. To keep representative samples of reoccurring classes, while gradually incorporating more data from recent classes, one can create a dual-memory FIFO mechanism that combines long-term inter-class balance with short-term intra-class responsiveness $c_{\text{txt}}=S_{t}\cup L_{t}$, as depicted in Figure 6. To maintain stable inter-class boundaries over time, a long-term memory $L_{t}\subset D_{<t}$ of size $M_{\text{long}}$ retains temporally older samples that ensure a balanced distribution across all known classes $\mathcal{Y}$. This memory helps preserve representation from rare or fading classes, especially when their frequency $p_{i}(y)\ll 1$. Over time, minority classes may evolve from outlier regions toward denser areas, or even swap roles with majority classes as the stream progresses. The long-term memory ensures that the context remains representative of the most up to date global distributional structure over extended periods. Complementarily, a short-term memory $S_{t}\subset D_{t-M_{\text{short}}:t}$ captures the most recent samples from the stream. This memory focuses on preserving intra-class variability and recent sub-distributions $p_{t}(x\mid y)$. Within a single class $y\in\mathcal{Y}$, examples may cluster into temporally local sub-concepts, separated in input space due to drift. These sub-clusters may expand, shift, or disappear with time. The short-term buffer tracks these fine-grained dynamics, enabling the model to adapt to class recurrence, sub-concept emergence, and local outlier patterns.
Figure 6. Short- and long-term memories
Table 1. Prequential accuracy (*: no statistical difference), NC = No-change, MC = Majority-Class
Dataset LTM ARF (n=90) BOLE (n=50) LevBag (n=45) SRP (n=80) EFDT VFDT NC MC
NOAA 81.29 ± 0.26 78.36 ± 0.20 73.45 ± 0.51 75.51 ± 0.23 78.35 ± 0.18 72.34 ± 0.43 70.86 ± 1.56 68.03 68.62
METER 73.62 ± 0.28 68.02 ± 0.36 67.28 ± 1.54 63.17 ± 5.78 69.69 ± 1.81 57.94 ± 2.21 52.07 ± 2.63 1.08 8.22
ELEC 91.82 ± 0.18 87.89 ± 0.22 90.17 ± 0.98* 85.86 ± 0.16 88.70 ± 0.57 77.96 ± 1.28 77.87 ± 1.29 85.30 57.41
RIALTO 79.03 ± 1.06 67.95 ± 0.36 47.27 ± 1.54 60.99 ± 5.78 75.12 ± 1.81 55.79 ± 2.21 28.68 ± 2.63 00.00 10.00
POSTURE 59.72 ± 0.18 59.14 ± 0.19* 48.28 ± 1.09 55.44 ± 0.69 58.14 ± 0.55 48.91 ± 0.18 52.99 ± 1.05 20.07 27.29
COVER 95.53 ± 0.23 91.75 ± 0.14 93.33 ± 0.59* 85.26 ± 0.24 94.54 ± 0.20* 82.38 ± 1.08 76.98 ± 2.37 95.06 48.76
POKER 97.51 ± 0.31 89.17 ± 0.43 80.01 ± 1.54 88.52 ± 3.12 87.58 ± 0.88 77.38 ± 4.10 77.43 ± 7.02 74.54 50.11
Average 82.65 77.47 71.40 73.54 78.87 67.53 62.41 49.15 38.63
Rank 1.00 2.86 4.57 4.29 2.71 6.00 6.57 7.43 8.43
FIFO algorithm. The pseudocode is given in Algorithm 1. It begins by initializing the short-term memory $S$ and long-term memory $L$, along with a class count dictionary $C[y]$ for each class $y$. The size of the long-term memory is computed as $M_{\text{long}}=M-M_{\text{short}}$. After a warming period $T_{\text{warm}}$, the algorithm enters the main loop, processing each new sample $(x_{t},y_{t})$ from the data stream $D_{\text{stream}}$. For each new sample, the algorithm first adds it to the short-term memory $S$. If the short-term memory exceeds its maximum size $M_{\text{short}}$, the oldest sample is removed and moved to the long-term memory $L$. The class count for the corresponding class of the removed sample is updated in the dictionary $C$. Once the long-term memory exceeds its capacity $M_{\text{long}}$, the algorithm identifies the most overrepresented class, $y_{\max}$, in long-term memory $L$. The oldest sample from this overrepresented class is then removed to maintain balance in the memory, ensuring that the system does not retain too many samples from any single class.
Algorithm 1 Naive context optimization
1:Input: $D_{\text{stream}}$, $M_{\text{short}}$, $M_{\text{long}}$, $T_{\text{warm}}$, LTMLTM